# Efficient Prior Calibration From Indirect Data

### v1

@@ 5 Numerical Results for Darcy Flow
Throughout we use regularization $h(\alpha)=\frac{1}{2\sigma_{h}^{2}}\|\log(\alpha)-m_{h}\|^{2}_{2}$ in which the log is applied component-wise; we refer to $\sigma_{h}$ and $m_{h}$ as the standard-deviation and mean of the regularizer term. For all experiments, we use 1000 slicing directions to evaluate the sliced-Wasserstein term, and we minimize the relevant loss functions using Adam [35] with a learning rate of $10^{-2}$ decayed by half four times on the parameters $\alpha$, other than where explicitly stated. All experiments are run on a 24GB NVIDIA RTX 4090 GPU.
@@ 5.1 Level Set Priors
Figure 1: (a) Spectrum decay of the square root of the eigen values for the covariance operator (20) (which corresponds to the standard deviaiton of the basis expansion coefficient of each mode $\varphi_{j}(x)$) for $\beta=4$ and different choices of lengthscale parameter $\lambda$. As $\lambda$ increases, more modes play a significant role. (b) A comparison of smooth level set function $\tilde{z}$ and sharp level set function $z$ transitioning at $x=0.5$ for $\tau=100$.
@@
We note that there is a symmetry in inference between $\kappa^{+}$ and $\kappa^{-}$, so after convergence we sort the $\kappa^{\pm}$ to associate $\kappa^{+}$ with the larger value and vice versa for $\kappa^{-}$. With this experimental setup established, we can test the proposed methodology. Subsections 5.1.1 and 5.1.2 concern the setting where we use a PDE solve for the forward model, and hence minimize Functional 2; we consider dimensions $d=1$ and $d=2$ respectively. In Subsections 5.1.3 and 5.1.4 we combine learning of $\alpha$ with operator learning to replace the PDE solve, and hence minimize Functional 3, Functional 3; we again consider dimensions $d=1$ and $d=2$ respectively.
@@ 5.1.1 Prior Calibration: 1D Darcy
We use $N=1000$ data $y$ each of $d_{y}=50$ noisy pointwise observations of the solution with $\Gamma=0.01^{2}\mathrm{I}$ observational noise covariance. The prior KL expansion is truncated at 20 terms. We use $N_{s}=1000$ samples to evaluate Functional 2. The true parameters for data generation are $\lambda=8,\kappa^{+}=2,\kappa^{-}=1$ and we set the level set smoothing parameter $\tau=10$. We use a finite element mesh of 100 nodes with a weak form residual computed through array shifting and solved with conjugate gradients and the forcing function $f$ is set to a constant value of 10 for all examples. No regularizer $h(\alpha)$ is used for the 1D example. We minimize using the smoothed model without sharp interfaces but we derive data both with the sharp and smoothed interfaces. In both cases we are able to accurately recover all three hyper parameters in $\alpha$: see Fig. 3. The runtime for these experiments is 41 seconds for the full 1k iterations. The relative error of the predicted parameters for the smoothed level set prior are $0.56\%$, $0.28\%$, and $0.96\%$ for $\kappa^{+}$, $\kappa^{-}$, and $\lambda$ respectively. We observe that the mis-specification between $\mu^{\dagger}$ (draws are discontinuous) and $(T^{\alpha})_{\#}\mu_{0}$ (draws are smooth) is reflected in a small inference bias, as would be expected. In Fig. 4 we plot the prior parameter estimation mean and standard deviations for 100 random initializations, varying the number of samples $N_{s}$ used for estimating the loss functions and the size of the dataset ($N$). The $\lambda$ parameters are initialized from $\log\lambda\sim\mathrm{Unif}(\log 0.5,\log 4)$, $\log\kappa^{-}\sim\mathrm{Unif}(\log 0.5,\log 4)$, and $\log\kappa^{-}\sim\mathrm{Unif}(\log 6,\log 10)$. We observe that the mean of the recovered parameters roughly converges for this problem setup after a dataset size $N=100$ and number of samples $N_{s}=100$.
@@ 5.1.2 Prior Calibration: 2D Darcy
In this section we generalize the setting of the previous section two the two-dimensional domain $D=[0,1]\times[0,1]$. As in the 1D example, we use $N=1000$ data $y$ each of $d_{y}=50$ noisy pointwise observations of the solution with $\Gamma=0.01^{2}\mathrm{I}$ observational noise covariance. The prior KL expansion is truncated at 20 terms in each dimension, for a total of 400 terms. We use $N_{s}=100$ samples to evaluate Functional 2. The true parameters for data generation are $\lambda=5,\kappa^{+}=2,\kappa^{-}=1$ and we set the level set smoothing parameter $\tau=5$. We use a finite element mesh of $100\times 100$ nodes. The regularizer $h(\alpha)$ in Functional 2 has means $m_{h,\lambda}=\log(10)$ and $m_{h,\kappa^{\pm}}=\log(3)$, with standard-deviations $\sigma_{h}=2$. In Fig. 5, two permeability fields associated with their sets of observations from the dataset are shown.
(a) $z$ field, 3
(b) Smoothed $z$ field, 3
(c) $u$ and 50 observations, 3
Figure 5: Plots of a sharp permeability fields and the smoothed field and PDE solutions with observation locations in the dataset for the 2D Darcy problems We not that the solutions are computed from the sharp level sets for the dataset, whereas $(T^{\alpha})_{\#}\mu_{0}$ in the loss function uses the smoothed level sets to make use of gradient-based optimization.
@@ 5.2.2 Unidentifiable Setting
For certain parameter values of $\alpha$, from which the true data is generated, there can be a lack of identifiability. In the previous section, the data was generated from a relatively rough random field with a relatively long length-scale. Hence the smoothness and length-scale parameters each have a distinct impact on the prior’s spectrum. We will now look at a smooth random field setting $\nu=4$, keeping $\ell=0.5$. Now, both parameters will have a similar influence on the spectrum of the prior. Other parameters for the data, solver/operator, and regularizer are kept the same as in Subsection 5.2.1. In Fig. 11 we show the convergence of the $\alpha$ parameters for the unidentifiable parameter regime for the $\alpha$ only learning and the joint $\alpha,\phi$ prior and operator learning tasks. We can see the lack of identifiability causes more issues in the operator learning as it struggles to gain any inferential traction. For these examples, it was necessary to run the learning models for more iterations (20k) and use a 6-time decaying learning rate on $\alpha$ (as opposed to the usual 4 in previous examples). Furthermore, due to training instability, the $\phi$ parameters had to be trained with AMSGrad [52](with the same settings) instead of Adam. As we can see from the Fig. 11, the proposed methodology applied to this particular unidentifiable setup does not fail silently, the $\alpha$ only learning has a very long convergence time, and the $\alpha$ and $\phi$ learning does not reach equilibrium. Other possible avenues to identifying non-uniqueness in prior parameter estimation include assessing the sensitivity of the estimated parameters on the choice of regularizer mean and standard-deviation, or the use of hierarchical parametric priors where a distribution on $\alpha$ is learned. There, a wide (or multimodal) inferred distribution on $\alpha$ could indicate non-identifiability of the prior parameters as well as the covariance between entangled prior parameters.
@@ 6 Conclusions
Funding
MG is grateful for support through EPSRC grants EP/X037770/1, EP/Y028805/1, EP/W005816/1, EP/V056522/1 , EP/V056441/1, EP/T000414/1, and EP/R034710/1. AMS is grateful for support through a Department of Defense Vannevar Bush Faculty Fellowship, from the Air Force Office of Scientific Research under MURI award number FA9550-20-1-0358 (Machine Learning and Physics-Based Modeling and Simulation) and by the SciAI Center, funded by the Office of Naval Research (ONR), under Grant Number N00014-23-1-2729. AV is supported through the EPSRC ROSEHIPS grant EP/W005816/1.
@@ Appendix A Proofs
Proof.
We first remember our definition
@@
By the definition of pushforward
(24) $\mathsf{W}^{2}_{2}(P_{\mathrm{B}\#}\nu,P_{\mathrm{B}\#}\mu)$ $=\inf_{\gamma\in\Pi(\nu,\mu)}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|\mathrm{B}^{-1/2}x-\mathrm{B}^{-1/2}y\|_{2}^{2}{\mathrm{d}}\gamma(x,y)$
(25) $=\inf_{\gamma\in\Pi(\nu,\mu)}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|x-y\|_{\mathrm{B}}^{2}{\mathrm{d}}\gamma(x,y).$

### v2

@@ 5 Numerical Results for Darcy Flow
For simple problems regularization may not be needed. For more challenging problems, regularization on $\alpha$ may help with numerical stability, particulariliy when the surrogate model $F^{\phi^{\star}(\alpha)}$ is concurrently learned. Hence, for the 1D Darcy problems we ommit the regularizer, and for the 2D Darcy examples we use regularization of the form $h(\alpha)=1/({2\sigma_{h}^{2}})\|\log(\alpha)-m_{h}\|^{2}_{2}$ in which the log is applied component-wise; we refer to $\sigma_{h}$ and $m_{h}$ as the standard-deviation and mean of the regularizer term. We have not found it necessary to regularize the neural operator parameters as we are not interested in recovering a specific value for $\phi^{\star}(\alpha)$, rather any value which achieves a small error when evaluating the PDE residual of the output is sought.
For all experiments, we use 1000 slicing directions, $\theta$, to evaluate the sliced-Wasserstein term, and we minimize the relevant loss functions using Adam [37] with a learning rate of $10^{-2}$ decayed by half four times on the parameters $\alpha$, other than where explicitly stated. All experiments are run on a 24GB NVIDIA RTX 4090 GPU.
@@ 5.1 Level Set Priors
Figure 1: (a) Spectrum decay of the square root of the eigen values for the covariance operator (20) (which corresponds to the standard deviation of the basis expansion coefficient of each mode $\varphi_{j}(x)$) for $\beta=4$ and different choices of lengthscale parameter $\lambda$. As $\lambda$ increases, more modes play a significant role. (b) A comparison of smooth level set function $\tilde{z}$ and sharp level set function $z$ transitioning at $x=0.5$ for $\tau=100$.
@@
We note that there is a symmetry in inference between $\kappa^{+}$ and $\kappa^{-}$, so after convergence we sort the $\kappa^{\pm}$ to associate $\kappa^{+}$ with the larger value and vice versa for $\kappa^{-}$. The number of physical systems from which we have data, $N$, is set to 1000 for the subsequent examples as this is shown to give very accurate results whilst highlighting the main benefit of the methodology – to efficiently and accurately estimate underlying distributional parameters from large amounts of data.
With this experimental setup established, we can test the proposed methodology. Subsections 5.1.1 and 5.1.2 concern the setting where we use a PDE solve for the forward model, and hence minimize Functional 2; we consider dimensions $d=1$ and $d=2$ respectively. In Subsections 5.1.3 and 5.1.4 we combine learning of $\alpha$ with operator learning to replace the PDE solve, and hence minimize Functional 3, Functional 3; we again consider dimensions $d=1$ and $d=2$ respectively.
@@ 5.1.1 Prior Calibration: 1D Darcy
We use $N=1000$ data $y$ each of $d_{y}=50$ noisy pointwise observations of the solution with $\Gamma=0.01^{2}\mathrm{I}$ observational noise covariance. The prior KL expansion is truncated at 20 terms. We use $N_{s}=1000$ samples to evaluate Functional 2. The true parameters for data generation are $\lambda=8,\kappa^{+}=2,\kappa^{-}=1$ and we set the level set smoothing parameter $\tau=10$. We use a finite element mesh of 100 nodes with a weak form residual computed through array shifting and solved with conjugate gradients and the forcing function $f$ is set to a constant value of 10 for all examples. No regularizer $h(\alpha)$ is used for the 1D example. We minimize using the smoothed model without sharp interfaces but we derive data both with the sharp and smoothed interfaces. In both cases we are able to accurately recover all three hyper parameters in $\alpha$: see Fig. 3. The runtime for these experiments is 41 seconds for the full 1k iterations. The relative error of the predicted parameters for the smoothed level set prior are $0.56\%$, $0.28\%$, and $0.96\%$ for $\kappa^{+}$, $\kappa^{-}$, and $\lambda$ respectively. We observe that the mis-specification between $\mu^{\dagger}$ (draws are discontinuous) and $(T^{\alpha})_{\#}\mu_{0}$ (draws are smooth) is reflected in a small inference bias, as would be expected. In Fig. 4 we plot the prior parameter estimation mean and standard deviations for 100 random initializations, varying the number of samples $N_{s}$ used for estimating the loss functions and the size of the dataset ($N$). The $\alpha$ parameters are initialized from $\log\lambda\sim\mathrm{Unif}(\log 0.5,\log 4)$, $\log\kappa^{-}\sim\mathrm{Unif}(\log 0.5,\log 4)$, and $\log\kappa^{-}\sim\mathrm{Unif}(\log 6,\log 10)$. We observe that the mean of the recovered parameters roughly converges for this problem setup after a dataset size $N=100$ and number of samples $N_{s}=100$.
@@ 5.1.2 Prior Calibration: 2D Darcy
In this section we generalize the setting of the previous section to the two-dimensional domain $D=[0,1]\times[0,1]$. As in the 1D example, we use $N=1000$ data $y$ each of $d_{y}=50$ noisy pointwise observations of the solution with $\Gamma=0.01^{2}\mathrm{I}$ observational noise covariance. The prior KL expansion is truncated at 20 terms in each dimension, for a total of 400 terms. We use $N_{s}=100$ samples to evaluate Functional 2. The true parameters for data generation are $\lambda=5,\kappa^{+}=2,\kappa^{-}=1$ and we set the level set smoothing parameter $\tau=5$. We use a finite element mesh of $100\times 100$ nodes. The regularizer $h(\alpha)$ in Functional 2 has means $m_{h,\lambda}=\log(10)$ and $m_{h,\kappa^{\pm}}=\log(3)$, with standard-deviations $\sigma_{h}=2$. In Fig. 5, two permeability fields associated with their sets of observations from the dataset are shown.
(a) $z$ field
(b) Smoothed $z$ field
(c) $u$ and 50 observations
Figure 5: Plots of a sharp permeability fields and the smoothed field and PDE solutions with observation locations in the dataset for the 2D Darcy problems We note that the solutions are computed from the sharp level sets for the dataset, whereas $(T^{\alpha})_{\#}\mu_{0}$ in the loss function uses the smoothed level sets to make use of gradient-based optimization.
@@ 5.2.2 Unidentifiable Setting
For certain parameter values of $\alpha$, from which the true data is generated, there can be a lack of identifiability. In the previous section, the data was generated from a relatively rough random field with a relatively long length-scale. Hence the smoothness and length-scale parameters each have a distinct impact on the prior’s spectrum. We will now look at a smooth random field setting $\nu=4$, keeping $\ell=0.5$. Now, both parameters will have a similar influence on the spectrum of the prior. Other parameters for the data, solver/operator, and regularizer are kept the same as in Subsection 5.2.1. In Fig. 11 we show the convergence of the $\alpha$ parameters for the unidentifiable parameter regime for the $\alpha$ only learning and the joint $\alpha,\phi$ prior and operator learning tasks. We can see the lack of identifiability causes more issues in the operator learning as it struggles to gain any inferential traction. For these examples, it was necessary to run the learning models for more iterations (20k) and use a 6-time decaying learning rate on $\alpha$ (as opposed to the usual 4 in previous examples). Furthermore, due to training instability, the $\phi$ parameters had to be trained with AMSGrad [57](with the same settings) instead of Adam. As we can see from the Fig. 11, the proposed methodology applied to this particular unidentifiable setup does not fail silently, the $\alpha$ only learning has a very long convergence time, and the $\alpha$ and $\phi$ learning does not reach equilibrium. Practitioners may also find the method is more sensitive to initializations in such unidentifiable settings. Other possible avenues to identifying non-uniqueness in prior parameter estimation include assessing the sensitivity of the estimated parameters on the choice of regularizer mean and standard-deviation, or the use of hierarchical parametric priors where a distribution on $\alpha$ is learned. There, a wide (or multimodal) inferred distribution on $\alpha$ could indicate non-identifiability of the prior parameters as well as the covariance between entangled prior parameters.
@@ Appendix A Proofs
Proof A.1.
Let $\nu,\mu$ denote two probability measures defined on a common measure space $(X,\Sigma).$ Let $\Pi(\nu,\mu)$ denote the set of all coupling probablity measures $\gamma$, on the product space $X\times X$, such that $\gamma(A\times X)=\nu(A)$, $\gamma(X\times A)=\mu(A)$ for all $A\in\Sigma$. Remembering $P_{\mathrm{B}}(\cdot)=\mathrm{B}^{-1/2}\,\cdot\,$, now recall the definition
@@
The set of all $\gamma^{\prime}\in\Pi(P_{\mathrm{B}\#}\nu,P_{\mathrm{B}\#}\mu)$ is equivalent to the set $(P_{\mathrm{B}}\otimes P_{\mathrm{B}})_{\#}\gamma$ defined over all $\gamma\in\Pi(\nu,\mu)$. Thus we have
(24) $\mathsf{W}^{2}_{2}(P_{\mathrm{B}\#}\nu,P_{\mathrm{B}\#}\mu)$ $=\inf_{\gamma^{\prime}\in\Pi(P_{\mathrm{B}\#}\nu,P_{\mathrm{B}\#}\mu)}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|x^{\prime}-y^{\prime}\|_{2}^{2}\,{\mathrm{d}}\gamma^{\prime}(x^{\prime},y^{\prime})$
$=\inf_{\gamma\in\Pi(\nu,\mu)}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|x-y\|_{2}^{2}\,{\mathrm{d}}((P_{\mathrm{B}}\otimes P_{\mathrm{B}})_{\#}\gamma(x,y))$
$=\inf_{\gamma\in\Pi(\nu,\mu)}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|\mathrm{B}^{-\frac{1}{2}}x-\mathrm{B}^{-\frac{1}{2}}y\|_{2}^{2}\,{\mathrm{d}}\gamma(x,y)$
$=\inf_{\gamma\in\Pi(\nu,\mu)}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|x-y\|_{\mathrm{B}}^{2}\,{\mathrm{d}}\gamma(x,y)$
(25) $=\mathsf{W}^{2}_{2,\mathrm{B}}(\nu,\mu).$